The tens digit of a two-digit number is 5 more than the units digit. If 3 is subtracted from the number and 2 is added to the reversed number, the former will be twice the latter. What is the number? Is it 2[10(x+5)+ x - 3] = 2 + 10(x)
+ x + 5? How is it written? Thank you.

Answer by scott8148(6628)

the number is 10t + u

t = u + 5

10t + u - 3 = 2(10u + t + 2) ___ 10t + u - 3 = 20u + 2t + 4 ___ 8t = 19u + 7

substituting ___ 8(u + 5) = 19u + 7 ___ 33 = 11u ___ 3 = u

substituting ___ t = (3) + 5 = 8

the number is 83